Find the equivalent capacitance.

C1=2.0 µF, C2=2.5 µF, C3=4.5 µF, C4=7.5 µF, C5=5.0 µF


1
Expert's answer
2021-01-06T14:00:38-0500

Parallel combination of Capacitor:-let all capacitor are connected in parallel combination,the equivalent capacitance


"C_{eq}=C_1+C_2+C_3+C_4+C_5\\C{eq}=2+2.5+4.5+7.5+5.0\\C{eq}=21.5\mu\digamma"

Series Combination of Capacitor:-let all capacitor are connected in series combination,the equivalent capacitance is


"\frac{1}{C_{eq}}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}+\frac{1}{C_4}+\frac{1}{C_5}\\\frac{1}{C_{eq}}=\frac{1}{2}+\frac{1}{2.5}+\frac{1}{4.5}+\frac{1}{7.5}+\frac{1}{5}\\\frac{1}{C_{eq}}=1.46\\C_{eq}=0.68\mu\digamma"